### v2

# Cross-lingual Contextualized Phrase Retrieval

Abstract
Phrase-level dense retrieval has shown many appealing characteristics in downstream NLP tasks by leveraging the fine-grained information that phrases offer. In our work, we propose a new task formulation of dense retrieval, cross-lingual contextualized phrase retrieval, which aims to augment cross-lingual applications by addressing polysemy using context information. However, the lack of specific training data and models are the primary challenges to achieve our goal. As a result, we extract pairs of cross-lingual phrases using word alignment information automatically induced from parallel sentences. Subsequently, we train our Cross-lingual Contextualized Phrase Retriever (CCPR) using contrastive learning, which encourages the hidden representations of phrases with similar contexts and semantics to align closely. Comprehensive experiments on both the cross-lingual phrase retrieval task and a downstream task, i.e, machine translation, demonstrate the effectiveness of CCPR. On the phrase retrieval task, CCPR surpasses baselines by a significant margin, achieving a top-1 accuracy that is at least 13 points higher. When utilizing CCPR to augment the large-language-model-based translator, it achieves average gains of 0.7 and 1.5 in BERTScore for translations from XEn and vice versa, respectively, on WMT16 dataset. We release our code and data at https://github.com/ghrua/ccpr_release.
Cross-lingual Contextualized Phrase Retrieval
Huayang Li♣††thanks: Work done during Huayang’s internship at Tencent AI Lab. Correspondence to Deng Cai and Lemao Liu. Deng Cai Zhi Qu♣ Qu Cui Hidetaka Kamigaito♣ Lemao Liu Taro Watanabe♣ ♣Nara Institute of Science and Technology ♡Tencent AI Lab {li.huayang.lh6, qu.zhi.pv5, kamigaito.h, taro}@is.naist.jp {jcykcai, qucui, redmondliu}@tencent.com
1 Introduction
Compared with the dense retrieval at sentence (or passage) level Karpukhin et al. (2020); Borgeaud et al. (2022); Asai et al. (2024), learning dense retrieval at the phrase level has shown more appealing characteristics in extensive NLP tasks, such as entity linking Gillick et al. (2019), open-domain question answering Lee et al. (2021a, b), text generation Lan et al. (2023); Cao et al. (2024), etc. An important reason is that phrases can provide more fine-grained information than sentences.
In cross-lingual research, the phrase-level dense retrieval also shows the promise to solve a range of NLP problems Bapna and Firat (2019); Zheng et al. (2022). The cross-lingual phrase retrieval in Zheng et al. (2022) focuses on the mapping of cross-lingual Wiki entities, where each of them is represented by averaged hidden representations from 32 different contexts. However, the constrained phrase type, i.e., the wiki entity, makes the cross-lingual phrase retriever difficult to augment general NLP tasks. Moreover, unlike wiki entities, which may have fewer ambiguities, general-type phrases that are lexically identical can have different meanings depending on their contexts. Thus, accounting for polysemy Cruse (1986) using the context information becomes critical. Even when the lexically identical phrases share similar semantics, a more appropriate context would provide richer information for utilizing the phrases in downstream tasks Min et al. (2019); Cao et al. (2024).
Therefore, we advocate for a new task formulation, i.e., cross-lingual contextualized phrase retrieval, which aims to find the cross-lingual phrase that is mostly relevant to the provided (general-type) source phrases, considering their meanings and surrounding contexts. However, achieving this goal is non-trivial, due to the scarcity of specific training data and models. Since annotating cross-lingual contextualized phrase pairs of general type is very difficult and expensive, we first introduce a data collection method, which leverages the automatically induced word alignment information from parallel sentences to extract suitable cross-lingual phrase pairs for training. This ensures that the phrases are of general type and both cross-lingual phrases and contexts are well aligned. Thereafter, we propose a Cross-lingual Contextualized Phrase Retriever (CCPR) which is trained on the constructed dataset. The CCPR mainly employs phrase-level contrastive learning to draw cross-lingual phrases with similar contexts and meanings closer in the hidden space. At inference time, we can leverage the well-trained CCPR to build a phrase-level index and use the query phases for search.
Unlike previous works Zheng et al. (2022), which focus solely on the task of cross-lingual phrase retrieval for wiki entities, our work also explores the potential of leveraging CCPR to augment downstream cross-lingual tasks, e.g., machine translation (MT). One critical question in front of us is how to select meaningful phrases for indexing and searching at inference time. To address this problem, we propose to learn a phrase segmentation module, which can be used to predict meaningful phrases from sentence- or passage-level retrieval data at inference time. The developed phrase segmentation module is important for ensuring the train-test consistency of CCPR.
Experiments on both of the cross-lingual contextualized phrase retrieval and a downstream task, i.e., MT, show the effectiveness of CCPR. We first evaluate CCPR on the cross-lingual contextualized phrase retrieval task, since a higher accuracy is more beneficial for downstream tasks, such as MT, cross-lingual dictionary induction, etc Zhang and Zong (2016); Søgaard et al. (2018). For this task, we use the human annotated cross-lingual phrase pairs as the golden truth and evaluate whether CCPR and other baselines can retrieve those golden phrases from a large-scale index. Experiments show that CCPR outperforms the baselines for at least 13 points in terms of the top-1 accuracy. We also conduct evaluation on the MT task using a large language model (LLM), e.g., LLama-2 Touvron et al. (2023). This task aims to evaluate whether the information retrieved by CCPR can enhance LLM’s ability of cross-lingual generation. Following the fashion of retrieval-augmented generation (RAG) Lewis et al. (2020b), we simply integrate the retrieved phrase information to the input of the LLM, and compare our method with other baselines. For both XEn and EnX translation directions on the WMT16 dataset, where X is from six languages, our CCPR achieves averaged gians of 0.7 and 1.5 BERTScore points Zhang et al. (2019), respectively. Results on other evaluation metrics, e.g., COMET Rei et al. (2022), are consistent with BERTScore.
In summary, our contributions are three-fold:
We propose a new formulation of dense retrieval, i.e., the cross-lingual contextualized phrase retrieval, which has substantial potential in augmenting cross-lingual tasks.
We propose a Cross-lingual Contextualized Phrase Retriever (CCPR), which uses the constructed training data, i.e., cross-lingual phrases extracted from automatically induced word alignment information, to learn both the phrase alignment and segmentation modules.
We conduct extensive experiments on cross-lingual contextualized phrase retrieval and MT. Our method outperforms the baselines by a large margin.
2 Related Work
Dense Retrieval
In text generation, dense retrieval has been widely studied at both sentence and phrase levels. Particularly after the emergence of large language models (LLMs), how to retrieve related sentences (or passages) Izacard et al. (2021); Ni et al. (2022) for input to LLMs has became a popular research topic, namely, retrieval-augmented generation (RAG) Lewis et al. (2020b); Karpukhin et al. (2020); Guu et al. (2020); Borgeaud et al. (2022); Asai et al. (2024). However, compared to the sentence-level retrieval, phrase-level retrieval has been shown to not only enhance retrieval accuracy but also markedly boost performance in downstream tasks Lee et al. (2021a, b). By indexing and integrating retrieved phrases into the output of language models, Lan et al. (2023) and Cao et al. (2024) make the text generation process more attributable and accurate. These studies underline the effectiveness of phrase-level dense retrieval. Our work notably diverges from those works, because we focus more on cross-lingual tasks.
Retrieval in Cross-lingual Tasks
In cross-lingual tasks, the majority of research has concentrated on sentence-level retrieval. Many research works in retrieval-augmented MT Zhang et al. (2018); Gu et al. (2018); Xia et al. (2019); He et al. (2021) have explored retrieving bilingual sentences based on similarity of source sentences, guiding MT models to utilize retrieved target information. A notable limitation is their reliance on bilingual data for constructing the index. This requirement substantially limits the scale and diversity of data available for retrieval. In contrast, Cai et al. (2021) propose to directly retrieve relevant target sentences using source sentences to assist MT, facilitating cross-lingual retrieval using a monolingual index. Unlike the task-specific model for MT in Cai et al. (2021), many studies in dense retrieval aim to enhance the cross-lingual sentence retrieval on more languages and at a larger scale Chidambaram et al. (2019); Conneau et al. (2020); Reimers and Gurevych (2020); Heffernan et al. (2022); Feng et al. (2022); Cai et al. (2022).
However, cross-lingual phrase retrieval has not been extensively investigated. Bapna and Firat (2019) utilize the source side of the bilingual data for building a retrieval index and employ the similarity between source phrases for retrieval. Therefore, their indexing also suffers from the limitation of reliance on bilingual data. In addition, it uses -grams as phrases by default, leading to plenty of meaningless and noisy spans in the index. Zheng et al. (2022) limit their cross-lingual phrase retrieval task to a specific phrase type, i.e., the wiki entity. The setting in their work overlooks the nuanced meanings that lexically identical phrases may convey in different contexts, making the retriever difficult to augment general NLP tasks. In contrast, our research aims to develop a cross-lingual contextualized phrase retriever for general-type phrases.
Additionally, there is research focused on employing nearest neighbor retrieval to support MT Khandelwal et al. (2021); Zheng et al. (2021); Meng et al. (2022); Deguchi et al. (2023). These efforts differ from the aforementioned studies as they aim to estimate prediction distributions through nearest neighbor retrieval at each translation step.
3 Task Formulation
This work aims a cross-lingual phrase retriever that can augment downstream NLP tasks. To ensure its effectiveness, the retriever must adhere to two essential criteria. First, it should ensure that phrases in the index are of general type. Second, it must resolve the polysemy of general-type phrases by considering their contextual differences.
Thus, we introduce a new task formulation, cross-lingual contextualized phrase retrieval. We are given a large collection of general-type phrases in language , and a general-type query phrase in language . Each phrase, by default, is associated with the context information , where represents the original sentence containing this phrase, and and are its start and end positions in . The primary objective is to identify a cross-lingual phrase that is relevant to , considering their contexts and meanings. Our formulation notably diverges from the one proposed by Zheng et al. (2022), which focuses on a restricted type of phrases, i.e., wiki entities, ignoring the nuanced meanings that general-type phrases may convey in different contexts.
This new task formulation also introduces several challenges, with the primary issue being the lack of training data. Specifically, there is a scarcity of cross-lingual phrase pairs of general type and accompanied by contextual information. Such data are crucial for training the model to recognize cross-lingual phrases that have similar contexts and meanings, allowing it to align these phrases closely in a hidden space. In addition, how to extract meaningful general-type phrases from sentences is also an open question, which is critical for employing the retriever to downstream tasks.
4 Training Data Collection
Annotating cross-lingual contextualized phrase pairs of general type is very difficult and expensive, posing a significant obstacle to training effective cross-lingual contextualized phrase retriever. However, the sentence-level parallel data in general domains are more readily available Ng et al. (2019). Additionally, as the lower unit of phrase alignment, the word alignment has been extensive studied Brown et al. (1993); Och and Ney (2003); Dyer et al. (2013); Li et al. (2019); Jalili Sabet et al. (2020); Dou and Neubig (2021); Wu et al. (2023). Therefore, to overcome the data scarcity, we propose to use a word alignment model, such as GIZA++ Och and Ney (2003) and neural aligner Dou and Neubig (2021), in order to automatically induce word alignments from parallel sentences and subsequently extract cross-lingual phrase pairs. This method allows us to produce phrase pairs that are not only of general type but also accompanied with contexts, aligning with our task requirements.
Formally, given a pair of parallel sentences , and , where and are the number of words in the sequences, we use a word alignment model to obtain the word alignment information of and Koehn et al. (2003); Och et al. (1999). If every word in a consecutive span can be aligned to a consecutive span , we will use them as a pair of aligned phrases and , where and . Notably, phrases that meet the constraints may range from a single word to the whole sequence, and extracted phrases are allowed to have overlaps. In addition, the context information of and are and , respectively. Unlike the phrase table extracted in SMT Koehn et al. (2003); Chiang (2005), which is static and context-independent, here the phrase mapping between and is unique considering their context.
5 Methodology
5.1 Model Architecture
We introduce a Cross-lingual Contextualized Phrase Retriever (CCPR) for our new formulated task. The main target of CCPR is to make the representations of cross-lingual phrases with similar contexts and meanings to be close in the hidden space. To this end, we propose a cross-lingual phrase alignment module based on contrastive learning, utilizing the data collected in Sec. 4 for training. In addition, one remaining problem is how to select phrases for indexing at inference time. To address this, we propose a phrase segmentation module that uses phrase representations to predict meaningful phrases from sentences or paragraphs. At inference time, we can leverage the learned phrase segmentation module to select phrases for indexing, ensuring the train-test consistency.
Phrase Encoding
Our phrase encoder is based on a context encoder and an layer. The encoding of a phrase , is defined as:
where is a Transformer model, e.g., BERT Devlin et al. (2019); Feng et al. (2022) and RoBERTa Conneau et al. (2020), that encodes the context to a matrix with a dropout mask . Inspired by preview works Lan et al. (2023); Lee et al. (2021a); Seo et al. (2018), we use the concatenation of and , i.e., the start and end token of the phrase, to represent phrase . is a function that maps the concatenation of two hidden states and from to , where is the output hidden size. The encoding for phrases in context is similar.
Cross-lingual Phrase Alignment
For a batch of parallel sentences, we can extract a collection of cross-lingual phrase pairs , where and are the positive examples of each other. Since each positive phrase pair is from parallel sentences that share similar semantics, directly learning on them may cause the model to learn a trivial shortcut. Therefore, inspired by SimCSE Gao et al. (2021), we apply two independently sampled dropout mask and , to encode phrases and , respectively, following Eq. 1. The training objective for the direction is:
where we use all the non-paired phrases in as in-batch negatives in Eq. 3. For bidirectional symmetry, our final loss is:
The illustration of the contrastive learning is shown in Fig. 1 (a). In practice, we may use parallel data from multiple languages during training, i.e., the source contexts in a batch may come from multiple languages, e.g., Germany, Czech, etc, which is inspired by the success of multi-lingual training in Liu et al. (2020).
Phrase Segmentation
This module aims to learn how to select phrases from sentences or passages, since most retrieval data at inference time comes in sentences or passages, rather than phrases. To ensure the train-test consistency, we take the phrases extracted in Sec. 4 as positive data, and all the other spans in the corresponding contexts as negative data. Our phrase segmentation module is a binary classifier defined as follows:
where is the label for the span , where , is an activation function, is linear layer that maps the hidden representation from to , and the definitions for and are the same as in Eq. 1. We use the label for phrases and for non-phrase spans. The phrase segmentation for sequence is similar to Eq. 5. In practice, the number of non-phrase spans is significantly more than the number of phrases. To mitigate data imbalance, we employ a strategy of randomly sampling an equal number of non-phrase spans and phrases within a sentence during training Li et al. (2020). The illustration of phrase segmentation is in Fig. 1 (b).
The training loss for the phrase segmentation is
where is a set of spans, which are extracted from sequence or . The learning objective of the whole model becomes:
where is a hyper-parameter. It is worth noting that the parameters of used for phrase alignment and segmentation are shared.
In our preliminary studies, we investigated various phrase segmentation strategies, including -gram segmentation and learning a Byte Pair Encoding (BPE) model Sennrich et al. (2016); Kudo (2018) across word boundaries. However, these methods struggle to identify meaningful phrases, leading to significant discrepancies between training and testing phases and performance declines. In Sec. 6.3, we evaluate the effect of the learned phrase segmentation on MT task.
5.2 Inference Pipeline
The pipeline of employing our cross-lingual contextualized phrase retriever for downstream NLP tasks involves two main steps: index building and searching, as shown in Fig. 1 (c).
A significant benefit of our retriever is its ability to leverage monolingual data in the target language, eliminating the need for bilingual data to construct the index at inference time Zhang et al. (2018); Gu et al. (2018); Bapna and Firat (2019). Given the retrieval data in target language, our retriever begins by segmenting sentences (or passages) into phrases. When segmenting a sentence (or passage), we will enumerate all the possible spans within it and calculate their probabilities of being meaningful phrases, according to Eq. 5. We select all the spans whose probabilities are larger than a threshold for indexing. It is worth noting that overlapping phrases are allowed at the phrase segmentation time. The selected phrases are encoded into hidden representations according to Eq. 1, which will be used to construct a dense-retrieval index.
For querying, we first process the query sentence (or paragraph) in a source language similarly, i.e., segmenting and encoding. Subsequently, we use the encoded query phrases to directly search for the related cross-lingual phrases using Maximum Inner Product Search (MIPS). Because of the advanced data structure and search algorithm Malkov and Yashunin (2018); Douze et al. (2024); Johnson (2022), the search step is highly efficient (more details in App. A.3). The retrieved information, including the target phrase, its accompanied context, and its positions in the context, can then be integrated to downstream tasks to improve the performance. As shown in Fig. 2, the retrieved phrases along with their surrounding contexts are integrated into the instructions given to an LLM for tasks such as MT.
| Model | DeEn | RoEn | CsEn | AVG. |
| XLMR | 1.5 | 0.0 | 4.0 | 1.8 |
| mBERT | 8.0 | 3.0 | 7.5 | 6.1 |
| mUSE | 34.5 | 30.0 | 45.5 | 36.6 |
| LaBSE | 52.0 | 33.5 | 61.0 | 48.8 |
[…]
| CPR-LaBSE | 57.0 | 60.0 | 78.5 | 65.1 |
[…]
| CCPR-LaBSE | 75.0 | 72.5 | 88.0 | 78.5 |
6 Experiments
We propose to evaluate our model on cross-lingual contextualized phrase retrieval and MT.
6.1 Implementation Details
Training Data Collection
We use all the bilingual training datasets of WMT16 on Huggingface111https://huggingface.co/datasets/wmt16 to train our Cross-lingual Contextualized Phrase Retriever (CCPR). The WMT16 dataset has six language pairs, inclduing De-En, Cs-En, Fi-En, Ru-En, and Tr-En. Because of the efficiency and the satisfactory performance of GIZA++ Och and Ney (2003); Dou and Neubig (2021), we first use the GIZA++ software222https://github.com/moses-smt/mgiza, i.e., IBM-4 model Brown et al. (1993), to induce the word alignment for each pair of parallel sentences, and then extract the cross-lingual phrase pairs. More details about phrase extraction are in App. A.1. The final training dataset contains 1.3 billion cross-lingual phrase pairs extracted from 10 million parallel sentences.
Model
In our work, we train two variants of our model, CCPR-XLMR and CCPR-LaBSE, whose parameters are initialized from XLM-RoBERTa-base (XLMR) Conneau et al. (2020) and LaBSE Feng et al. (2022), respectively. The hidden size of both models is 768. The output of in Eq. (1) is 128, which plays a critical role in reducing the memory cost of our phrase index while maintaining comparable performance. At inference time, we use the FAISS library333https://github.com/facebookresearch/faiss Douze et al. (2024) to build our retrieval index. We use the FlatIP as our index type. Training details are App. in A.2
| Model (XEn) | DeEn | CsEn | FiEn | RuEn | RoEn | TrEn | AVG. |
| Llama-2-7B | 75.42 | 69.03 | 66.13 | 71.65 | 72.89 | 55.10 | 68.37 |
| + XLMR | 75.32 | 69.15 | 66.06 | 71.74 | 73.07 | 55.49 | 68.47 |
| + LaBSE | 75.74 | 69.60 | 66.81 | 72.12 | 73.22 | 56.76 | 69.04 |
[…]
| + CCPR-LaBSE | 75.96 | 70.05 | 67.17 | 72.28 | 73.57 | 59.22 | 69.70 |
| Model (EnX) | EnDe | EnCs | EnFi | EnRu | EnRo | EnTr | AVG. |
| Llama-2-7B | 63.42 | 52.32 | 48.96 | 84.18 | 82.99 | 58.47 | 65.05 |
| + XLMR | 64.03 | 52.34 | 49.04 | 84.08 | 82.99 | 58.66 | 65.19 |
| + LaBSE | 64.26 | 52.90 | 49.76 | 84.28 | 83.05 | 60.02 | 65.71 |
[…]
| + CCPR-LaBSE | 64.29 | 54.93 | 51.54 | 84.97 | 83.36 | 64.07 | 67.19 |
6.2 Cross-lingual Contextualized Phrase Retrieval
As outlined in Sec. 3, given a query phrase and its context in source language, the objective of this task is to identify the most relevant cross-lingual phrase from a large-scale index in target language, considering the context information.
Setup
We build the test set based on the human annotated word alignment data Jalili Sabet et al. (2020). We process the word alignment data on three language pairs, DeEn Ghader and Monz (2017), RoEn Mihalcea and Pedersen (2003), and CsEn Mareček (2011). For the CsEn dataset, we only leverage the data under the “pcedt” split, which are data from The Wall Street Journal (WSJ). For each language pair, we instruct human annotators to identify 200 high-quality phrase pairs. The details of the human annotation process are discussed in App. B.1. These annotated source phrases serve as queries, while their aligned target phrases help construct the index. To mimic real-world conditions, where an index contains extra data, we use 9.6 million English phrases sampled from the WMT16 training dataset as the extra data.
We compare our models, CCPR-XLMR and CCPR-LaBSE, with several baselines. The first two baselines XLMR Conneau et al. (2020) & mBERT Devlin et al. (2019) 444https://huggingface.co/models are models trained on multilingual data. We use them to encode the context and represent the phrases by concatenating the hidden representations of their start and end tokens. The second two baselines mUSE Chidambaram et al. (2019) & LaBSE Feng et al. (2022) are trained on cross-lingual data, and we use the model weights released by SBERT555https://www.sbert.net Reimers and Gurevych (2020) to encode the contexts. To evaluate the effect of context-awareness, we propose two context-independent counterparts, i.e., CPR-XLMR and CPR-LaBSE. Compared with our contextualized models, the only difference is that the CPR-X methods use cross-lingual phrase pairs in semantically different contexts as positive examples, similar to the setup in Zheng et al. (2022). We build an independent index for each model. In this task, we use the top-1 accuracy to evaluate our methods and baselines.
| Model | Index | EnRo | EnCs | EnTr | AVG. |
| CCPR-XLMR | Train | 82.89 | 53.66 | 62.03 | 66.19 |
| CCPR-XLMR | NC | 83.18 | 54.53 | 63.40 | 67.03 |
[…]
| CCPR-LaBSE | NC | 83.36 | 54.93 | 64.07 | 67.45 |
Results
Tab. 1 shows that our models significantly outperform baselines in cross-lingual contextualized phrase retrieval, highlighting the capability of our models to accurately identify relevant cross-lingual phrases while being context-aware. In contrast, the CPR-LaBSE and CPR-XLMR, i.e., the context-independent baselines inspired by Zheng et al. (2022), perform notably worse.
6.3 Machine Translation
The evaluation on MT task is to assess if the information retrieved by our models can enhance performance on downstream tasks. Rather than training a new MT model Vaswani et al. (2017), we propose to integrate the information to the prompt of the LLM, e.g., Llama-2 Touvron et al. (2023), following the fashion of retrieval-augmented generation (RAG) Lewis et al. (2020a). Thus, we can fairly evaluate the effect of the retrieved information.
Setup
We choose Llama-2-7B Touvron et al. (2023) model as the backbone model for MT. We use two of our phrase retrieval models for this task, i.e., CCPR-Labse and CCPR-XLMR, as discussed in Sec. 6.1. We also consider a multilingual and a cross-lingual sentence retrieval models as baseline methods, i.e., the XLM-RoBERTa-base (XLMR) Conneau et al. (2020) and LaBSE Feng et al. (2022). The two sentence-level baselines are proposed to evaluate whether the phrase-level retrieval information is more beneficial for augmenting LLM’s cross-lingual ability.
We integrate the retrieval information into input of the LLM. An example of the prompt of our method is shown in Fig. 2, where the retrieved target phrases with their contexts are presented along with the source phrases in the prompt. The presented contexts are truncated to no more than 100 characters in our method. More details about prompts of our method and other baselines are shown in App. C.1. For our methods, when building the index, we first use the learned segmentation module (Sec. 5.1) to extract contextualized phrases from the retrieval data and then encode them into dense vectors using our models. For the XLMR and LaBSE baselines, we directly use them to encode sentences in retrieval data and build a sentence-level index. For all methods, we use the monolingual newscrawl data of 2016 666https://data.statmt.org/news-crawl/${LANG}/news.2016.${LANG}.shuffled.deduped.gz as the retrieval data. More setup details are in App. C.
Main Results
As shown in Tab. 2, both of our CCPR-XLMR and CCPR-LaBSE outperform the baseline methods when assisting the LLM with the information from cross-lingual contextualized phrase retrieval. Especially on some low-resource directions, such as the TrEn, EnCs, EnFi, CCPR-LaBSE outperforms other baselines by more or around 2 BERTScore points. This indicates that the the contextualized cross-lingual phrase information is more critical for improving the performance on low resource languages, because the percentage of Cs, Fi, and Tr data in Llama-2 training are all less than Touvron et al. (2023). Consistent results on COMET Rei et al. (2022) are shown in App. C.4.
In addition, we find a good correlation between the performance of our cross-lingual contextualized phrase retrieval task and the MT task, showing that higher retrieval accuracy leads to superior results in downstream applications.
| Segmentation | EnRo | EnCs | EnTr | AVG. |
| N-gram | 83.16 | 53.42 | 60.74 | 65.77 |
| PSM | 83.36 | 54.93 | 64.07 | 67.45 |
Analysis
In Tab. 2, all retrieval-based methods construct their index using the monolingual newscrawl data, which is approximately six times the size of the bilingual training data. It highlights a key benefit of our cross-lingual phrase retrieval: the ability to utilize extensive monolingual resources to build the index Cai et al. (2021). Therefore, in Tab. 3, we evaluate the differences of indexing on monolingual data and the bilingual data. We observe that exploring the vast monolingual data leads to significantly better performance. In addition, although the -gram is widely used for phrase segmentation in previous works Bapna and Firat (2019); Lee et al. (2021a, b); Lan et al. (2023), our experiments demonstrate that a learned phase segmentation is more suitable for augmenting cross-lingual tasks, as shown in Tab. 4.
7 Conclusion & Future Works
This paper introduces a novel approach to dense retrieval, i.e., cross-lingual contextualized phrase retrieval, focusing on resolving phrase polysemy by utilizing contextual information. The main challenge identified is the scarcity of training data, specifically cross-lingual phrase pairs with context. To overcome this, we use a word alignment model to derive such phrase pairs from parallel sentences. We then present the Cross-lingual Contextualized Phrase Retriever (CCPR), which employs contrastive learning to effectively capture similar meanings and contexts of cross-lingual phrases. Our extensive testing across retrieval and machine translation tasks shows that CCPR significantly outperforms existing baselines.
8 Limitations
While our cross-lingual contextualized phrase retrieval holds substantial potential, it is not without limitations. Notably, the phrase-level index required by our approach is considerably larger than that of a sentence-level index, given that each sentence can encompass numerous phrases. This expansion necessitates increased disk space for storing the index, requiring additional engineering techniques to maintain the scalability of our cross-lingual contextualized phrase retriever (CCPR), e.g., index quantization or sharding. Furthermore, to improve the performance of large language models (LLMs) on cross-lingual tasks, it becomes necessary to integrate the information from multiple retrieved phrases of the query sentence into the LLM input. This integration process can lead to a rise in the inference costs associated with LLMs.
9 Ethical Statement
We focus on leveraging cross-lingual contextualized phrase retrieval to augment the performance of downstream NLP tasks. We emphasize that our model is strictly designed and applied in a manner that avoids the generation of sensitive information, such as disinformation or content aimed at deceiving individuals. Furthermore, we assure that all data utilized for the training and evaluation of our model have been sourced from publicly accessible datasets. Our commitment to ethical research ensures our work benefits the NLP field responsibly, without compromising the ethical standards.
Appendix A Implementation Details
A.1 Extraction of Cross-lingual Phrase Pairs
Because of the efficiency and the satisfactory performance of GIZA++ Och and Ney (2003); Dou and Neubig (2021), we first use the GIZA++ software777https://github.com/moses-smt/mgiza, i.e., IBM-4 model Brown et al. (1993), to induce the word alignment for each pair of parallel sentences, and then extract the cross-lingual phrase pairs. Notably, we did not induce the phrase table as in Phrase-based Statistical Machine Translation (PBSMT) Koehn et al. (2003). Instead, in our setting, we extract the cross-lingual consecutive spans that are aligned in a pair of parallel sentence as cross-lingual phrase pairs. Each of the cross-lingual phrase pairs is associated with the surrounding contexts. We filter out the phrases which begin or end with words whose frequency is more than 30k in the training dataset. We also filter out the phrases that contain only numbers and punctuations.
A.2 Training of CCPR
Each model is trained on the mixture of all language pairs of WMT16. We train our model on 8 V100 GPUs for 20K steps, where the learning rate is 5e-5, dropout rate is 0.2, batch size on each device is 64. The for the phrase segmentation loss in Eq. 5 is . Those hyper-parameters are tuned on the validation dataset.
A.3 Inference Latency
In our experiments of cross-lingual contextualized phrase retrieval, retrieving the top-32 nearest neighbors for 1000 source phrases took only 0.04 seconds on 4 V100 GPUs, thanks to FAISS’s high parallelism on GPUs Douze et al. (2024). We believe this retrieval latency is adequate for most real-time retrieval scenarios.
Appendix B Cross-lingual Contextualized Phrase Retrieval
B.1 Human Annotation
We asked human annotators to label 200 high-quality phrase pairs as the golden truth data for each of the language pairs, i.e., DeEn, RoEn, and CsEn. Hiring human annotators to annotate those bilingual data is expensive and time consuming. However, fortunately, some bilingual datasets already provide the human-annotated word alignment Ghader and Monz (2017); Mihalcea and Pedersen (2003); Mareček (2011). Therefore, in our task, we only ask the human annotators, three authors of our work, to annotate the English side of the data, i.e., whether a English span is a high-quality and meaningful phrase or not.
More concretely, we first use heuristic rules as discussed in Sec. 6.1 to collect an initial set of cross-lingual phrase pairs. For each phrase pair, we ask the human annotator to answer three questions:
1.
If the phrase is a single word, whether it is informative in the context?
2.
If the phrase has multiple words, whether the semantics of this phrase is complete and informative? For instance, “local authorities and large” is not a phrase with complete semantics. In addition, “Of course” has complete semantics but is not informative.
If the answer of any questions is true, then we add this phrase to a pool of high-quality phrases. For those phrases, overlaps are allowed. We finally randomly sampled 200 annotated phrases form the pool for each language.
Appendix C Machine Translation
C.1 Translation Instructions
C.2 Building Index
| EnDe | EnCs | EnRo | |
| CoT | 55.20 | 47.02 | 80.22 |
| CCPR-LaBSE | 64.29 | 54.93 | 83.36 |
The indexing strategies diverge between the baseline methods and our approach. For sentence-level baselines, i.e., XLMR and LaBSE, the indexing is straightforward, directly using the sentences. In contrast, our method employs a learned phrase segmentation module to extract phrases from sentences for indexing. When applying our model to MT, we set the phrase segmentation thresholds in Eq. 5 to for indexing and for querying. The rationale behind a lower indexing threshold is to populate the index with a broader array of phrases, enhancing its comprehensiveness. Conversely, a higher threshold for query documents ensures that the segmented phrases are more accurate. These hyper-parameters were tuned using the validation dataset from WMT16.
C.3 Format Alignment
Since the output format of the Llama-base model is hard to control and the Llama-chat model, i.e., the model after instruction-tuning, may reject to translate a sentence that contains negative words, we decide to fine-tune the Llama base model using LoRA Hu et al. (2022) for 100 steps to align the translation format. We use Platypus888https://github.com/arielnlee/Platypus Lee et al. (2023) as our code base. We follow all the hyper-parameters of Platypus to fine-tune the LoRA module. The data we used for fine-tuning contain 1,800 parallel sentences randomly sampled from the WMT16 training set. For each method, we only fine-tune one model for the XEn (or EnX) direction, where X is the non-English language.
C.4 More Results
More results on evaluated by COMET Rei et al. (2020) are shown in Tab. 5, which is consistent with the results evaluated by BERTScore Zhang et al. (2019).
Additional, to evaluate that the effectiveness of the retrieved information, we also compare with a Chain-of-Though (COT) approach Wei et al. (2022). As shown in Figure 8, simply translating phrases before sentence translation does not achieve the expected performance gain and is slightly worse than the baseline approach. In contrast, our approach outperforms this method by a large margin. One hypothesis is that our approach is able to leverage external knowledge, which is more helpful for the translation task.